### v1

# Bridging the Rossby number gap in rapidly rotating thermal convection

Bridging the Rossby number gap in rapidly rotating thermal convection
Adrian van Kan,1† Keith Julien,2111Deceased on 14th of April 2024.   Benjamin Miquel,3 Edgar Knobloch1
1 Department of Physics, University of California at Berkeley, Berkeley, CA 94720, USA
2 Department of Applied Mathematics, University of Colorado, Boulder, CO 80309, USA
3 CNRS, École Centrale de Lyon, INSA Lyon, Université Claude Bernard Lyon 1,
Laboratoire de Mécanique des Fluides et d’Acoustique, UMR5509, F-69134 Écully, France
To whom correspondence should be addressed; E-mail: avankan@berkeley.edu
Geophysical and astrophysical fluid flows are typically buoyantly driven and are strongly constrained by planetary rotation at large scales. Rapidly rotating Rayleigh-Bénard convection (RRRBC) provides a paradigm for direct numerical simulations (DNS) and laboratory studies of such flows, but the accessible parameter space remains restricted to moderately fast rotation (Ekman numbers ${\rm Ek}\gtrsim 10^{-8}$), while realistic ${\rm Ek}$ for astro-/geophysical applications are significantly smaller. Reduced equations of motion, the non-hydrostatic quasi-geostrophic equations describing the leading-order behavior in the limit of rapid rotation (${\rm Ek}\to 0$) cannot capture finite rotation effects, leaving the physically most relevant part of parameter space with small but finite ${\rm Ek}$ currently inaccessible. Here, we introduce the rescaled incompressible Navier-Stokes equations (RiNSE) – a reformulation of the Navier-Stokes-Boussinesq equations informed by the scalings valid for ${\rm Ek}\to 0$. We provide the first full DNS of RRRBC at unprecedented rotation strengths down to $\rm Ek=10^{-15}$ and below, and show that the RiNSE converge to the asymptotically reduced equations.
Introduction
The universe abounds with examples of turbulent flows that are driven by buoyancy and constrained by rotation. These highly complex flows shape our environment across scales: in protoplanetary disks, they mediate the formation of planetesimals  (?); in planetary atmospheres, including on Earth, they carry heat, momentum and moisture, crucially impacting the climate  (?, ?, ?); in the Earth’s oceans  (?, ?), they transport heat, carbon dioxide, and nutrients, thereby crucially influencing biotopes  (?); in planetary and stellar cores, they are responsible for the generation of large-scale magnetic fields via the dynamo instability  (?, ?, ?); in the subsurface oceans of the gaseous giants’ moons, they shape the icy crusts of Europa, Enceladus, Titan, etc.  (?, ?, ?, ?); in stars like the Sun  (?, ?) they are responsible for large-scale magnetic fields as well as rich dynamics such as the 22-year cycle in solar activity.
Such geo- and astrophysical flows, in spite of their differences, typically share two important properties. Owing to their turbulent nature, their dynamics result from a broad range of interacting spatial and temporal scales. Accordingly, advection by the flow dominates over viscous forces when the Reynolds number is large, ${\rm Re}_{H}=UH/\nu\gg 1$, based here on the fluid layer depth $H$, a typical velocity of the flow $U$, and kinematic viscosity $\nu$. Simultaneously, advection by the flow occurs on time scales much longer than the planetary rotation period, as measured by small Rossby numbers ${\rm Ro}_{H}=U/(2\Omega H)\ll 1$, where $\Omega$ is the planetary rotation rate.    Also of particular importance is the ratio of the viscous and Coriolis forces providing an a priori external parameter referred to as the Ekman number ${\rm Ek}={\rm Ro}_{H}/{\rm Re}_{H}=\nu/(2\Omega H^{2})$, as well as the Prandtl number $\sigma=\nu/\kappa$ characterizing the fluid under consideration, where $\kappa$ is the thermal diffusivity. The strength of the thermal forcing is controlled by the Rayleigh number ${\rm Ra}$, which is proportional to the temperature drop $\Delta_{T}$ across the fluid layer (see Methods and Materials section).
Celestial body ${\rm Ek}$ $\rm Pr$ ${\rm Ro}_{H}$ ${\rm Re}_{H}$
Earth’s Outer Core $10^{-15}$ $0.1$ $10^{-7}$ $10^{8}$
Mercury (core) $10^{-12}$ $0.1$ $10^{-4}$ $10^{8}$
Jupiter (core) $10^{-19}$ $0.1$ $10^{-10}$ $10^{9}$
     Europa (ocean) $10^{-12}$ $11.0$ $10^{-2.5}$$10^{-1.5}$ $10^{9.5}$$10^{10.5}$
     Ganymede (ocean) $10^{-10}$$10^{-13}$ $10.0$ $10^{-3.5}$$10^{1.5}$ $10^{9.5}$$10^{11.5}$
Saturn (core) $10^{-18}$ 0.1 $10^{-9}$ $10^{9}$
      Enceladus (ocean) $10^{-10}$$10^{-11}$ $13.0$ $10^{-3.5}$$10^{-1}$ $10^{7.5}$$10^{9}$
     Titan (ocean) $10^{-11}$$10^{-12}$ $10.0$ $10^{-3}$$1$ $10^{9}$$10^{11}$
Neptune (core) $10^{-16}$ 10.0 $10^{-6}$ $10^{10}$
Uranus (core) $10^{-16}$ 10.0 $10^{-6}$ $10^{10}$
Sun (convection zone) $10^{-15}$ $10^{-6}$ $10^{-3}$ $10^{12}$
Table 1: Nondimensional parameter estimates for planetary  (?), satellite  (?) and stellar interiors  (?, ?). Estimates of the Rossby number are derived from the relation $\rm Ro_{H}=Re_{H}Ek$.
Celestial bodies typically feature extreme values for the nondimensional parameters defined above, as summarized in Table 1: rapid planetary rotation is reflected in very low Ekman and Rossby numbers, whereas large Reynolds numbers indicate the highly turbulent character of the flow. Such rapidly rotating, highly turbulent flows are challenging to analyze and specifically to simulate numerically owing to the sheer number of spatial and temporal degrees of freedom which they involve. One common approach for obtaining theoretical predictions at geophysically relevant parameters (specifically, ${\rm Ek}\rightarrow 0)$ relies upon identifying transitions between different flow regimes in a region of more moderate parameters (${\rm Ek}\gtrsim 10^{-8}$) that is amenable to direct numerical simulations (DNS) or laboratory experiments, together with the scaling laws for transport coefficients valid within each regime. These results are then extrapolated to extreme parameters on the assumption that the regimes and scaling laws observed for accessible parameter values extend to geo-/astrophysically relevant parameter values. Assuming that no nontrivial transition occurs outside the observed moderate parameter interval is a particularly strong assumption, akin to a leap of faith due to the current paucity of data at the most extreme parameter values.
The dynamics of the fluid flows found in the celestial bodies listed in Table 1 are further rendered complex by additional ingredients such as spherical geometry, multiple contributors to the density, compressibility, and the presence of magnetic fields. Absent such complexities, the quintessential paradigm for investigating rotationally influenced buoyant flows is provided by rapidly rotating Rayleigh-Bénard convection (RRRBC). A large number of studies have been published on this model system, which is very well suited for detailed experimental, numerical and theoretical studies, see for example  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and the recent review  (?). In RRRBC, gravity is often taken to be antiparallel to the rotation on the assumption that the Froude number ${\rm Fr}_{\Omega}\equiv\Omega^{2}H/g$ is small. The resulting configuration is translation-invariant and believed to be relevant to the polar regions of planets, moons and stars as well as to their interiors, although the case of misaligned gravity and rotation axis has also been studied  (?, ?, ?, ?, ?).
A major obstacle in the study of rapidly rotating convection, even in the highly idealized setting of RRRBC, is the difficulty, to date, of attaining the extremely low Ekman numbers, typically $10^{-12}$ or smaller (cf. Table 1), that characterize geo- and astrophysical flows. By contrast, current state-of-the-art DNS of the full Navier-Stokes-Boussinesq equations that describe RRRBC have been restricted to ${\rm Ek}\gtrsim 10^{-8}$  (?, ?, ?, ?), despite enormous computational efforts. Inclusion of magnetic effects in the simulations to capture the capability for dynamo action, as well as spherical shell geometry, further limits the reported range of nearly all studies to ${\rm Ek}\gtrsim 10^{-7}$, ${\rm Re}_{H}\lesssim 10^{3}$  (?, ?, ?, ?, ?), although $\rm Ek=10^{-8}$ has been achieved in this context (?) but at great computational expense; simulations down to ${\rm Ek}=10^{-10}$ are available for linearized dynamics only  (?). In state-of-the-art laboratory experiments, the accessible parameter space is even more restricted, to ${\rm Ek}\gtrsim 10^{-7}$ (?, ?, ?, ?, ?). These accessible parameter values in numerical or laboratory experiments are orders of magnitude larger than the estimates for realistic geo-/astrophysical applications. On the other hand, existing reduced equations  (?), known as the non-hydrostatic quasi-geostrophic (NHQG) equations, describe the dynamics in the limit ${\rm Ek}\to 0$, but their applicability is by construction confined to asymptotically small values of ${\rm Ek}$. This leaves an unexplored gap corresponding to small but finite Ekman numbers, precisely the regime that is most relevant for geophysical and astrophysical applications.
It is extremely rare to have a numerical or laboratory approach that allows one to span the parameter range from prototypical laboratory and DNS settings all the way to realistic geophysical and astrophysical settings. Remarkably, taking advantage of the characteristics of rapidly rotating thermal convection leads to a new numerical approach that relies on appropriately rescaled equations, and enables exploration of this very wide parameter range.
Results
Below, we introduce the rescaled Navier-Stokes equations (RiNSE), an equivalent reformulation of the full Navier-Stokes-Boussinesq equations, informed by the scaling behavior valid in the limit of vanishing ${\rm Ek}$  (?, ?). The RiNSE formulation is described in detail in the Materials and Methods section, and in another paper  (?), where it is validated against existing results in the literature, and the full details of the numerical implementation are given. In short, the RiNSE retain all terms in the Navier-Stokes-Boussinesq equations while rescaling each variable according to its leading-order asymptotic scaling behavior valid for ${\rm Ek}\to 0$ in terms of the natural small parameter of rapidly rotating convection $\varepsilon\equiv{\rm Ek}^{1/3}$, see  (?, ?). In particular, the rapid rotation generates a high degree of anisotropy, which is explicitly incorporated in the RiNSE formulation: vertical length scales are nondimensionalized by the layer height $H$, while horizontal length scales are significantly smaller, given by $\varepsilon H$. Consequently, time is measured in units of the viscous diffusion time based upon $\varepsilon H$. Further details are given in the Materials and Methods section. As shown in  (?), this procedure ensures an adequate preconditioning of the equations of motion at small but finite ${\rm Ek}$, thus facilitating DNS in this highly relevant and yet unexplored regime.
Here, we present the first DNS of RRRBC at realistically small Ekman numbers for planetary, satellite and stellar interiors, using the RiNSE formulation, and analyze the physics of the flow in this extreme parameter range. Our extensive set of simulations was performed using the pseudospectral solver Coral  (?) for a periodic domain of horizontal extent $10\ell_{c}$, where $\ell_{c}\approx 4.82{\rm Ek}^{1/3}$ is the critical wavelength in the limit ${\rm Ek}\to 0$  (?). Specifically, we consider a fluid with $\sigma=1$ subject to stress-free boundaries (for simplicity) at top and bottom, where constant temperatures are imposed, and periodic boundary conditions in the horizontal directions. We initialize our simulations with small-amplitude noise about the rest state, except when continuing simulations at higher resolution. We explore unprecedentedly small but finite Ekman numbers, as low as ${\rm Ek}=10^{-15}$, values that are more than six orders of magnitude smaller than those previously attainable in DNS or in the laboratory.
Figure 1 shows the parameter space explored in our simulations. The top panel shows the estimates of these parameters for planetary interiors and oceans on icy satellites given in Table 1 in the plane spanned by the Ekman number ${\rm Ek}$ and the bulk Reynolds number ${\rm Re}_{H}$, together with blue circles showing the values reached in our simulations, indicating that we attain realistically small ${\rm Ek}$ equal to the estimated values for the outer core of the Earth (${\rm Ek}\approx 10^{-15}$) and the convective zone of the Sun (not shown, cf. Table 1). We further indicate the approximate range of parameters reached in previous numerical and laboratory studies, which are restricted to higher ${\rm Ek}$. Dashed lines highlight important regime boundaries where the Rossby number ${\rm Ro}$ equals unity in different parts of the flow. The green dashed line is an estimate of where the bulk of the flow is expected to lose rotational constraint, namely ${\rm Ro}_{H}=1$, i.e., ${\rm Re}_{H}=1/{\rm Ek}$. By contrast, the red dashed line shows the theoretical prediction for the transitional (convective) Rossby number ${\rm Ro}_{c}\equiv{\rm Ek}\sqrt{\mathrm{Ra/Pr}}$ where rotational constraint is lost in the boundary layer  (?), i.e., $\rm{\rm Ro}_{c,BL}={\rm Ek}^{1/5}$. This prediction was derived by Julien et al.  (?) in the low $\rm Ek$ limit, and has thus far been impossible to test in DNS owing to the prohibitive difficulty of simulating flows in this regime.
In rapidly rotating flows, higher values of $\rm Ra$ are required to trigger convection as a consequence of the Taylor-Proudman constraint (?). Specifically, the critical Rayleigh number for convection in the limit of small ${\rm Ek}$ is proportional to ${\rm Ek}^{-4/3}$ as shown by Chandrasekhar  (?). Therefore, it is natural to introduce $\rm\widetilde{Ra}\equiv RaEk^{4/3}$, the reduced Rayleigh number, which measures the supercriticality of rapidly rotating convection. The lower panel of Fig. 1 shows an alternative representation of the parameter space in terms of the reduced Rayleigh number $\rm\widetilde{Ra}$ and the Taylor number ${\rm Ta}={\rm Ek}^{-2}$. We reach Taylor numbers up to $10^{30}$ and reduced Rayleigh numbers up to $300$, corresponding to rapidly rotating and highly turbulent flows. The associated bare Rayleigh number $\rm Ra$, indicated by colors in the background, ranges over a similarly wide interval from $O(10^{3})$ to very large values in excess of $10^{20}$.
This wide array of numerical simulations allows us to probe, for the first time in DNS, the physics of rotating convection from moderate Ekman numbers down to the fully rotationally constrained regime, previously only accessible in asymptotically reduced equations  (?, ?) within the $\rm Ek\to 0$ limit.
[…]
Figure 1: Overview of the parameter space. Panel A: Estimates of the nondimensional parameters for different celestial objects (see Table 1) shown in the plane spanned by the Ekman and Reynolds numbers, with ${{\rm Ro}_{H}}={{\rm Re}_{H}}{\rm Ek}$. The blue circles indicate the parameter values explored here. Parameters reached in previous laboratory experiments and simulations are indicated by shaded regions and dashed lines indicate regime boundaries discussed in the text. Panel B: Overview of three different sets of simulations (${\rm\widetilde{Ra}=}const.$ shown by white diamonds, ${\rm Ra}=const.$ shown by black triangles and ${\rm Ek}=const.$ shown by blue plus signs) in the plane spanned by the reduced Rayleigh number $\rm\widetilde{Ra}=Ra\,Ek^{4/3}$ and the Taylor number $\rm Ta=Ek^{-2}$. Dashed lines represent the same regime boundaries as in the top panel.
Bulk Flow Morphology
[…]
Figure 2: Top row: vertically averaged (i.e., barotropic) vertical vorticity field $\overline{\omega_{z}}(x,y)$ at ${\rm Ek}=10^{-15}$ (panel A), $10^{-8}$ (panel B), $10^{-7}$ (panel C), $10^{-6}$ (panel D) and $\rm\widetilde{Ra}=120$ (all panels). Panel E: skewness of the barotropic vorticity field versus $\rm Ek$ indicating a strong cyclone-anticyclone asymmetry for $\rm Ek\gtrsim 10^{-9}$, and an approximate (statistical) cyclone-anticyclone symmetry for $\rm Ek\lesssim 10^{-9}$.
An important feature of rapidly rotating flows is the formation of large-scale flow structures, associated with spectral condensation of kinetic energy at the domain scale as a result of an inverse energy cascade  (?, ?, ?). This behavior is also observed more broadly in anisotropic turbulent flows, such as stably stratified rotating turbulence, turbulence subject to strong magnetic fields or geometric confinement to a thin-layer geometry  (?, ?). In the case of a periodic, horizontally square Cartesian domain, in the limit of asymptotically large rotation ${\rm Ek}\to 0$ the condensate takes the form of a large-scale vortex dipole, observed using reduced equations (?, ?), resembling findings in two-dimensional turbulence (?). By contrast, direct numerical simulations of rotating convection at previously attainable moderately small Ekman and Rossby numbers reveal an asymmetry between cyclones and anticyclones, namely a preference for strong cyclonic structures and diffuse, weaker anticyclones  (?). While a dominance of cyclones over anticyclones is also observed in stably stratified flows at moderately small Rossby numbers  (?, ?), the transition from that regime at moderately small Rossby numbers to very small ${\rm Ek}$ (relevant to the celestial bodies listed in Table 1), where the cyclone-anticyclone symmetry is expected to be restored, could not previously be reached in RRRBC. To address this question, we illustrate and analyze in Fig. 2 how the vertically averaged (i.e., barotropic) vertical vorticity $\overline{\omega_{z}}(x,y)$ (see the Materials and Methods section) changes as ${\rm Ek}$ varies for a fixed supercriticality $\rm\widetilde{Ra}=120$. At moderate rotation rates ($\rm{\rm Ek}=10^{-6}$ and ${\rm Ro}\approx 0.4$, panel D), a striking asymmetry exists between the strong, coherent cyclone and a weak, incoherent anticyclonic region. Here the a posteriori Rossby number $\rm Ro$ is computed as the root mean square of the ratio $\left\|\mathbf{u}\cdot\nabla\mathbf{u}_{\perp}\right\|/(2\Omega\left\|\mathbf{u}_{\perp}\right\|)$, where ${\mathbf{u}}_{\perp}=(u,v,0)$. As rotation increases (top row of Fig. 2, right to left), this asymmetry decreases but remains visible for ${\rm Ek}=10^{-8}$ (${\rm Ro}\approx 0.07$, panel B), a value close to the current state of the art for RRRBC studies. Finally, the cyclone-anticyclone symmetry expected in the asymptotic rapidly rotating limit ${\rm Ek}\to 0$ based on the NHQG equations is in fact observed for ${\rm Ek}=10^{-15}$ (${\rm Ro}\approx 8\times 10^{-5}$, panel A). To quantify this qualitative observation, we display in Fig. 2E the skewness of the barotropic vertical vorticity: $\mathrm{Skew}(\overline{\omega_{z}})$. A finite positive value, indicating the prevalence of strong cyclones, is obtained for ${\rm Ek}\gtrsim 10^{-9}$, whereas an approximately vanishing skewness is observed at smaller Ekman numbers, indicating a restored (statistical) cyclone-anticyclone symmetry, with a relatively sharp transition between the two regimes.
Boundary Layer Flow Morphology
[…]
Figure 3: Panel A: Overview of boundary layer flow morphology in terms of vertical vorticity $\omega_{z}$, temperature perturbation $\theta$ and vertical velocity $w$ at different ${\rm Ek}$ and $\rm\widetilde{Ra}=120$. Panel B: Histograms of $\omega_{z},\theta,w$ at different ${\rm Ek}$. Near-Gaussian statistics are observed at ${\rm Ek}=10^{-15}$ with larger skewness at larger ${\rm Ek}$. Panel C: Skewness of $\omega_{z},\theta,w$ versus ${\rm Ek}$. All results shown are obtained at $z=\delta_{\omega_{z}}$ (top of the momentum boundary layer near the bottom boundary), cf. Fig 5.
In addition to the barotropic component of the flow, which informs us about the bulk flow structure, the flow inside the thermal boundary layer near the stress-free, constant-temperature boundaries at the top and bottom is of great importance. Figure 3A shows visualizations of the flow inside the bottom thermal boundary layer (at height $z=\delta_{\omega_{z}}$, defined later, cf. Fig. 5) at $\rm\widetilde{Ra}=120$ for four distinct Ekman numbers ${\rm Ek}=10^{-15},10^{-8},10^{-7},10^{-6}$ in terms of three fields: the vertical vorticity $\omega_{z}$, the temperature perturbation $\theta$ away from the horizontally averaged temperature profile and the vertical velocity $w$ at the top of the (bottom) boundary layer, where the flow transitions between qualitatively distinct regimes. Panel B shows histograms of each field at the same Ekman numbers and panel C displays the associated statistical skewness versus ${\rm Ek}$. At very low Ekman numbers, $\rm Ek=10^{-15}$, where the bulk flow features two large-scale vortices and displays an approximate statistical cyclone-anticyclone symmetry, the same symmetry is obeyed in the boundary layer as evidenced by the approximately vanishing skewness not only of $\omega_{z}$, but also $\theta$ and $w$ (see also Supplementary Movies 1-3). A near perfect correlation can be seen between the vertical vorticity and $\theta$. We note that the RiNSE formulation remains numerically stable even when the Ekman number is reduced further, as far as $\rm Ek=10^{-24}$; the corresponding flow fields obtained from DNS are shown in the Supplementary Material. As the Ekman number is increased, the approximate symmetry of the histograms persists for a range up to $\rm Ek\approx 10^{-9}$ (not shown). At $\rm Ek\approx 10^{-8}$, where cyclone-anticyclone asymmetry and cyclone dominance emerges in the barotropic mode, the boundary layer flow morphology is also drastically altered: in addition to the trace of the strong large-scale cyclone, similarly strong, albeit short-lived, anticyclonic vortices emerge in the boundary layer, whose anticyclonic core is seen in some cases to be surrounded by a cyclonic shield (see also Supplementary Movie 4). These strong vortices with anticyclonic cores have a clear signature in the temperature fluctuation $\theta$ as strong, cold perturbations (in the bottom boundary layer shown here, see Supplementary Movie 5), and also in the vertical velocity $w$ which shows descending flow inside the vortex core, surrounded by rising fluid motion (see Supplementary Movie 6). As a result, the statistics of $\theta$ and $w$ are skewed towards negative values at this $\rm Ek$. It is important to stress that the structures displayed in Fig. 3 and discussed above are confined within the boundary layer, and do not extend into the bulk flow. To the best of our knowledge, the strong, partially shielded, short-lived anticyclonic vortices in the boundary layer with stress-free boundaries are described here for the first time, although shielded structures have been previously observed in various studies of rotating convection  (?, ?, ?, ?). We also note that similar structures were predicted based on Ekman layer theory for no-slip boundary conditions in  (?) in response to a Gaussian temperature perturbation. As ${\rm Ek}$ is increased to $10^{-7}$ and $10^{-6}$, cyclones clearly dominate the flow in the boundary layer (as in the bulk), while the strong localized anticyclones become more diffuse, leading to a significant skewness in the statistics of $\omega_{z}$. The responses in $\theta$ and $w$ remain visible in the boundary layer, particularly outside the large-scale cyclone, although their skewness decreases with further increase in $\rm Ek$.
[…]
Figure 4: Panel A: Sample time series of $\mathrm{Nu}-1$ at $\rm Ek=10^{-10}$ for ${\rm\widetilde{Ra}}=40,60,80$ in the statistically stationary state, revealing turbulent fluctuations about a well-defined mean value. Both mean and amplitude of fluctuations increase with $\rm\widetilde{Ra}$. Panel B: Same data, collapsed by rescaling the $y$ axis by ${\rm\widetilde{Ra}}^{3/2}$. Panel C: Temporally averaged $\mathrm{Nu}-1$ in the stationary state, compensated by ${\rm\widetilde{Ra}}^{3/2}$, with error bars indicating standard deviation, vs $\mathrm{Ta}={\rm Ek}^{-2}$. $\rm Nu$ increases with $\rm Ta$ at fixed $\rm\widetilde{Ra}$ up to a maximum, beyond which it decreases to converge to a $\rm Ta$-independent constant, which agrees with the value obtained in the NHQG equations at the same $\rm\widetilde{Ra}$ (horizontal dashed line), within one standard deviation (shaded area). Vertical dashed lines (same color scheme) correspond to $\mathrm{Ro_{c}}=1$ and $\mathrm{Ro}_{BL}=1$. Panel D: Corresponding data for the Reynolds number $\rm Re$, compensated by $\rm\widetilde{Ra}$, show a similar structure with an overshoot and eventual convergence to the NHQG prediction at large $\rm Ta$. Panel E: a posteriori Rossby numbers versus ${\rm Ek}^{1/3}$ for the bulk and the boundary layer, both of which exhibit collapse in panel F when shown versus ${\rm Ek}^{1/3}{\rm\widetilde{Ra}}^{5/4}$.
[…]
Averaging over time series like those shown in the top row of Fig 4, we obtain the Nusselt and Reynolds number statistics in panels C and D of Fig. 4, shown versus the Taylor number ${\rm Ta}={\rm Ek}^{-2}$. Specifically, four sets of simulations at constant $\rm\widetilde{Ra}=40,60,80,120$ are shown for a wide range of $10^{2}\leq\rm Ta\leq 10^{30}$. Triangles in different orientations for each $\rm\widetilde{Ra}$ show the observed average $\rm Nu$ and $\rm Re$, while error bars indicate the measured standard deviation about those mean values. Horizontal dashed lines indicate, for each $\rm\widetilde{Ra}$, the values corresponding to the reduced NHQG equations, with a shaded region indicating the corresponding standard deviation measured in the NHQG equations. At small $\rm Ta$ (large $\rm Ek$), for the given $\rm\widetilde{Ra}$, the flow is only weakly supercritical and therefore features only small $\rm Nu$ and $\rm Re$. A break is seen in the curves around ${\rm Ta}=10^{6}$ (${\rm Ek}=10^{-3}$), where the bulk convective Rossby number equals one, marked by faint vertical dashed lines, indicating that the bulk flow begins to be rotationally constrained. The inverse cascade sets in around $\rm Ta=10^{8}$ for $\rm\widetilde{Ra}=40$, and closer to $\rm Ta=10^{10}$ for larger $\rm\widetilde{Ra}=60,80,120$ (hollow circular markers). This is in agreement with the literature, see e.g.  (?). As $\rm Ta$ is increased (${\rm Ek}$ decreased) further from this point, an overshoot in $\rm Nu$ is observed leading to a maximum in the convective heat transport for the given control parameter $\rm\widetilde{Ra}$, whose amplitude and location depend on $\rm\widetilde{Ra}$. This is followed by convergence within one standard deviation to the value obtained in the NHQG equations at yet larger $\rm Ta$, beyond a threshold indicated by a second set of vertical dashed lines, one for each $\rm\widetilde{Ra}$, corresponding to convective Rossby number in the boundary layer of order unity, a condition equivalent to ${\rm Ek}={\rm\widetilde{Ra}}^{-15/4}$  (?), i.e., the rotation rate where the boundary layer also becomes rotationally constrained. These regime boundaries are equivalent to those shown in the top panel of Fig. 1. A similar structure is observed in the Reynolds number data, which collapse slightly less satisfactorily under the rescaling (this is known to be a consequence of the presence of the large-scale vortex  (?)). At small $\rm Ta$, the flow is only weakly supercritical and therefore the Reynolds number is small. It increases with $\rm Ta$ up to a maximum, beyond which it decreases to converge (within one standard deviation) to the value obtained from the NHQG equations. Since Taylor numbers in most works in the literature using laboratory experiments or unrescaled DNS have been limited to ${\rm Ta}\lesssim 10^{14}-10^{16}$, this $\rm Nu$ and $\rm Re$ vs $\rm Ta$ dependence has not previously been described. In particular, the presence of a maximum in the heat transport and turbulence intensity at a finite Taylor (or equivalently, Ekman) number as well as the statistical convergence to the NHQG equations are reported here for the first time.
In addition to the a priori (bulk/boundary layer) convective Rossby numbers shown by the dashed vertical lines in panels C and D of Fig. 4, panels E and F show the a posteriori Rossby number $\rm Ro$ computed as the root-mean-square of the ratio $\left\|\mathbf{u}\cdot\nabla\mathbf{u}_{\perp}\right\|/(2\Omega\left\|\mathbf{u}_{\perp}\right\|)$ in the bulk and in the boundary layer, where it takes distinct values (see Supplementary Materials). When shown versus $\varepsilon={\rm Ek}^{1/3}$, the standard small parameter of rotating convection, the data while scattered clearly show that the bulk Rossby number increases monotonically with ${\rm Ek}$ as expected, crossing unity near the observed onset of the inverse cascade. It is interesting that this occurs at Ekman numbers which are two orders of magnitude smaller than what may naively be expected based on the convective Rossby number $\rm Ro_{c}=1$. When the same data is plotted against the rescaled parameter $\rm Ek^{1/3}\widetilde{Ra}^{5/4}$, it collapses to a remarkable degree, both in the bulk and in the boundary layer. We stress that this is somewhat surprising, since this effective small parameter is obtained explicitly from the derivation of the reduced equations for the boundary layer  (?, ?). However, we note that the bulk flow develops cyclone-anticyclone symmetry at similar Ekman numbers as the boundary layer, which is another empirical link between the bulk and boundary layer flows observed here. The collapse of the data in Fig. 4E,F implies in particular that when $\rm Ek^{1/3}\widetilde{Ra}^{5/4}=1$, below which the average Nusselt and Reynolds numbers converge to the NHQG data, there is a well-defined associated critical Rossby number in the boundary layer, which is found to be approximately $0.2$. This is where the boundary layer loses rotational constraint and becomes unstable to strong ageostrophic motions with pronounced vertical gradients. These additional vertical gradients are associated with additional dissipation contributions. In steady state, the dissipation rates of kinetic energy and temperature variance are directly linked to the Nusselt number, via the so-called power integrals. In the RiNSE formulation, these read
$\langle|\nabla_{\perp}\theta|^{2}\rangle+\varepsilon^{2}\langle(\partial_{z}\theta)^{2}\rangle+\langle(\partial_{z}\overline{\Theta})^{2}\rangle=$ ${\rm Nu}-1,$ (1)
$\langle|\partial_{x}\mathbf{u}|^{2}+|\partial_{y}\mathbf{u}|^{2}\rangle+\varepsilon^{2}\langle|\partial_{z}\mathbf{u}|^{2}\rangle=$ $\frac{\rm\widetilde{Ra}}{\sigma^{2}}(\rm Nu-1),$ (2)
[…]
From the above analysis, it is apparent that the boundary layers play a nontrivial role in the transition from small to large ${\rm Ek}$, controlling the departure from the NHQG branch. To further characterize the boundary layer structure, we measure and analyze vertical root-mean-square profiles of different quantities as a function of $z$. Figure 5A shows a log-log plot of the thermal boundary layer depth $\delta_{\theta}$ computed as the location of the maximum in the root-mean-square vertical profile of $\theta$ (shown in the inset for different ${\rm Ek}$ at $\rm\widetilde{Ra}=120$). At small Ekman numbers, $\delta_{\theta}$ converges to a constant value for a given reduced Rayleigh number, corresponding to the NHQG limit. Panel C confirms that the thermal boundary layer depth scales with $\rm\widetilde{Ra}^{-15/8}$, which is the scaling obtained from reduced NHQG equations for the boundary layer  (?, ?). As ${\rm Ek}$ is increased close to the point where the boundary layer loses rotational support (${\rm Ek}=\rm\widetilde{Ra}^{-15/4}$, not shown), the boundary layer depth undershoots before increasing with a power law not far from $\rm Ek^{1/2}$ (and steeper than ${\rm Ek}^{1/3}$, not shown). Figure 5B shows an alternative measure of the boundary layer thickness, based on the root-mean-square vertical profile of $\partial_{z}\omega_{z}$. The inset shows sample profiles at $\mathrm{\widetilde{Ra}}=120$ for different Ekman numbers. A local maximum is seen in these profiles, which we use to define a momentum boundary layer thickness $\delta_{\omega_{z}}$. Since we are considering stress-free boundaries, where no linear Ekman layer is present, the existence of this momentum boundary layer is nontrivial. We interpret $\delta_{\omega_{z}}$ as a measure of how far the effect of the stress-free boundary penetrates into the bulk of the flow.
[…]
Figure 5: Boundary layer structure for a range of $\rm Ek$ and $\rm\widetilde{Ra}$. Panel A: log-log plot of the nondimensional width of the thermal boundary layer $\delta_{\theta}/H$, given by the location in $z$ of the maximum in the vertical r.m.s. profiles of $\theta$ shown in the inset, versus ${\rm Ek}$ for different $\rm\widetilde{Ra}$. Dashed line indicates a $\rm Ek^{1/2}$ power law. Panel B: log-log plot of the momentum boundary layer thickness $\delta_{\omega_{z}}$based on the location of the maxima of the r.m.s. profile of $\partial_{z}\omega_{z}$ (shown in the inset). Panel C: log-log plot of the nondimensional thermal boundary layer thickness $\delta_{\theta}/H$ at $\rm Ek=10^{-15}$ compared with the power law $\rm\widetilde{Ra}^{-15/8}$ predicted by the NHQG equations  (?).
We have furthermore verified that the momentum boundary layer seen here in the RiNSE formulation is also present in the reduced NHQG equations (not shown), although to our knowledge this has not been described previously. Like $\delta_{\theta}$, at sufficiently small ${\rm Ek}$, $\delta_{\omega_{z}}$ becomes constant, while at larger values of ${\rm Ek}$, there is again an undershoot followed by an power-law scaling regime where $\delta_{\omega_{z}}$ increases approximately as ${\rm Ek}^{1/2}$. The approximate ${\rm Ek}^{1/2}$ scaling, which is observed both in terms of $\delta_{\theta}$ and in $\delta_{\omega_{z}}$ and is also observed for no-slip boundary conditions  (?), is surprising in the present case, given that we consider stress-free boundaries, which do not lead to the formation of linear Ekman layers. While a boundary layer thickness proportional to ${\rm Ek}^{1/2}$ is theoretically predicted for the transitional thermal boundary layer  (?), the presence of such a scaling at significantly larger Ekman numbers, where the boundary layer flow is no longer rotationally dominated, goes beyond this prediction and the precise origin of this scaling law remains to be elucidated.
The results presented thus far were obtained along a particular cut through the physical control parameter space, namely varying $\rm Ek$ at fixed $\widetilde{Ra}$. This is the natural approach in the framework of the RiNSE. By contrast, keeping $\widetilde{Ra}$ fixed in an experimental setup or in an unrescaled DNS requires tuning two parameters. In such contexts, it is instead more natural to vary $\rm Ra$ at fixed $\rm Ek$ (equivalent to varying $\rm\widetilde{Ra}$ at fixed $\rm Ek$) or vary $\rm Ek$ at fixed $\rm Ra$ (equivalent to varying $\rm\widetilde{Ra}$ at fixed $\rm Ra$). To make contact with existing experimental and numerical studies, the corresponding cuts through parameter space are shown in Fig. 6. Panel A shows $\rm Nu-1$ versus $\rm\widetilde{Ra}$ for different sets of RiNSE simulations, each with $\rm Ra$ fixed to a different value between $\rm Ra=10^{6}$ and $\rm Ra=10^{10}$, as well as the corresponding values from the NHQG equations ($\rm Ra\to\infty$) at the same $\rm\widetilde{Ra}$. The NHQG data approximately follow the turbulent $\rm\widetilde{Ra}^{3/2}$ scaling law, cf. Fig. 4, while at finite $\rm Ra$, there is a well-defined transitional value $\rm\widetilde{Ra}_{t}$ of $\rm\widetilde{Ra}$ where $\rm Nu-1$ intersects the $\rm Ra\to\infty$ limit, and subsequently follows a shallower slope. This structure of the Nusselt number flattening out beyond a threshold $\rm\widetilde{Ra}=\widetilde{Ra}_{t}$ is robust among the datasets of runs with different $\rm Ra$. Qualitatively similar behavior was recently reported for RRRBC with no-slip boundaries  (?), although the eventual low ${\rm Ek}$ regime remains elusive in that case. Panel B shows the corresponding Reynolds number data for the same simulations. As discussed earlier, the corresponding turbulent scaling law $\rm Re\propto\widetilde{Ra}$ has proved to be more elusive owing to the impact of the inverse cascade, see  (?) and Fig. 4. Panel B shows a similar structure to panel A, in that the $\rm Re$ curves intersect the curve corresponding to the NHQG equations at a (numerically slightly different) threshold value for each $\rm\widetilde{Ra}$. Since these runs are performed at constant $\rm Ra$, each $\rm\widetilde{Ra}$ is equivalent to an Ekman number $\rm Ek$. Therefore $\rm\mathrm{Ra}_{t}$ is equivalent to a transitional Ekman number $\rm Ek_{t}$.
[…]
Figure 6: Nusselt number (panel A) and Reynolds number (panel B) versus the reduced Rayleigh number $\rm\widetilde{Ra}$ at fixed bare Rayleigh number $\rm Ra$. Dashed lines indicate intersection of a given ${\rm Ra}=const.$ set with the asymptotic limit. Transitional Ekman number (panel C) and reduced Rayleigh number (panel D) vs. $\rm\widetilde{Ra}$ at fixed bare Rayleigh number $\rm Ra$. $\rm Nu-1$ (panel E) and $\rm Re$ (panel F) versus $\rm\widetilde{Ra}$ at fixed Ekman number.
Panel C shows $\rm Ek_{t}$, corresponding to the vertical dashed lines in panels A and B, indicating that the data is quantitatively compatible with a scaling law $\rm Ek_{t}\propto Ra^{-5/8}$. This scaling corresponds to the theoretical prediction  (?) for the Ekman number at which the boundary layer loses rotational constraint. Here, we recover this scaling from a simple measurement of the Nusselt and Reynolds numbers. An equivalent representation of the same data is shown in panel D in terms of the reduced Rayleigh number $\rm\widetilde{Ra}$ versus the bare Rayleigh number $\rm Ra$, which also reveals satisfactory agreement with the corresponding theoretical scaling prediction $\widetilde{\rm Ra}_{t}\propto\rm Ra^{1/6}$.
Panels E and F of Fig. 6 show a third possible cut through the parameter space of rotating convection (qualitatively resembling results presented in  (?)), holding $\rm Ek$ fixed and changing $\rm Ra$ (or equivalently changing $\rm\widetilde{Ra}$). The turbulent scaling laws with $\rm\widetilde{Ra}$ compare similarly well with these datasets and also exhibit similar changes in scaling in the Nusselt and Reynolds numbers with $\rm\widetilde{Ra}$ at finite $\rm Ek$, up to ${\rm Ek}=10^{-6}$, while for ${\rm Ek}=10^{-9}$, ${\rm Ek}=10^{-10}$, the values of $\rm\widetilde{Ra}$ attainable in our simulations were not sufficiently high to observe a departure from the curve corresponding to the NHQG equations.
The flattening of the Nusselt number curves with increasing $\rm\widetilde{Ra}$ at finite Ekman numbers is a robust signal that could potentially be reproduced in laboratory experiments. Our results in Fig. 6 qualitatively resemble similar findings recently obtained for RRRBC with no-slip boundaries  (?).
Discussion
The results described above show that the RiNSE formulation, based on low Ekman number asymptotics, allows efficient direct numerical simulations far beyond the current state of art in laboratory experiments or unrescaled DNS, down to ${\rm Ek}=10^{-15}$ and below, which is comparable to the estimated values of ${\rm Ek}$ in the outer core of the Earth and the convection zone of the Sun. Specifically, this indicates that the RiNSE enable DNS at Ekman numbers which are over six orders of magnitude smaller than the smallest value obtained previously, in an impressive effort and at great computational cost, by means of unrescaled DNS  (?, ?). Using the RiNSE formulation, we revealed rich new physics in this previously inaccessible parameter regime. First, we uncovered a transition at ${\rm Ek}\approx\rm\widetilde{Ra}^{-15/4}$ ($\rm Ek\approx 10^{-8}$ for $\rm\widetilde{Ra}=120$) from a regime at larger ${\rm Ek}$ where the depth-averaged flow features a strong large-scale cyclone and a weak diffuse anticyclone towards a regime characterized by cyclone-anticyclone symmetry and a large-scale vortex dipole, in agreement with the predictions of the NHQG equations. Second, we identified a non-trivial transition in the boundary layer dynamics corresponding to the loss of rotational support in the boundary layer (confirming a previously untestable theoretical prediction of  (?)), a transition accompanied by the emergence of strong, albeit short-lived, anticyclonic structures near the ${\rm Ek}$ threshold that weaken as ${\rm Ek}$ increases, leading to dominance of strong cyclonic structures. It is interesting to note that the cyclone-anticyclone symmetry-breaking in the bulk approximately coincides in $\rm Ek$ with the boundary layer transition, although a theoretical explanation of this observation remains unavailable. While flows in the Earth’s atmosphere are characterized by only moderately small Rossby numbers and are therefore dominated by strong cyclones, our results suggest that in the interior of the Earth or other celestial bodies, cyclones and anticyclones may be statistically of equal strength.
We showed quantitatively that the time-averaged Nusselt and Reynolds numbers in steady state reflect these regime transitions, taking values consistent with the NHQG limit for ${\rm Ek}\lesssim\rm\widetilde{Ra}^{-15/4}$, while overshooting as ${\rm Ek}$ is increased at fixed $\rm\widetilde{Ra}$, reaching maximum heat transport near the value of ${\rm Ek}$ where the bulk Rossby number reaches unity, and decreasing to close to zero as ${\rm Ek}$ increases further at fixed $\rm\widetilde{Ra}$ due to reduced supercriticality. The increase in the Nusselt number is shown to be tied to increased dissipation associated with a change in the boundary layer flow morphology. The boundary depth was quantified in terms of the thermal fluctuations and vertical vorticity, revealing that at ${\rm Ek}\lesssim\rm\widetilde{Ra}^{-15/4}$ the boundary layer depth becomes ${\rm Ek}$-independent, taking the value associated with the NHQG equations, but undershooting for ${\rm Ek}\gtrsim\rm\widetilde{Ra}^{-15/4}$ before increasing approximately with an Ekman-like scaling $\rm Ek^{1/2}$ as ${\rm Ek}$ increases. Owing to the presence of stress-free boundary conditions, this behavior cannot be explained by a linear Ekman layer and is therefore a nonlinear effect. Finally, we have considered alternative cuts through the parameter space, with one set of runs varying ${\rm Ek}$ at fixed $\rm Ra$ and another set of runs varying $\rm Ra$ at fixed ${\rm Ek}$. This procedure revealed that at finite Ekman numbers, the Nusselt and Reynolds numbers remain close to the values found in the NHQG equations as $\rm\widetilde{Ra}$ increases, but start to deviate from them at a value of $\rm\widetilde{Ra}$ that depends on $\rm Ra$ in a way consistent with boundary layer theory  (?). We mention that the NHQG equations apply to systems with no-slip boundaries since in the limit ${\rm Ek}\to 0$ the no-slip boundaries become effectively stress-free  (?). At finite ${\rm Ek}$, however, Ekman boundary layers and the associated Ekman pumping necessarily modify the boundary layer structure, an effect that can be included in the NHQG equations following  (?).
The results presented here are based on the RiNSE reformulation of the equations and explain why previous state-of-the art DNS and laboratory experiments could not reach the parameter regime for observing the transition to fully rotationally constrained dynamics. The RiNSE formulation therefore opens the door to the further exploration of the parameter regime of very small but finite Ekman and Rossby numbers, highly relevant to planetary, satellite and stellar interiors. This is all the more timely since current and future observational missions prompted by the last Planetary Decadal Survey  (?), including the recently launched ESA’s JUICE (JUpiter Icy Moons Explorer) mission  (?) and NASA’s Europa clipper mission  (?) will provide new data requiring interpretation based on faithful model simulations at parameters as close as possible to realistic values. Our results are a first step in this direction.
RRRBC as studied here, with antiparallel gravity and rotation axis is an appropriate, albeit idealized model of the North pole regions of the celestial bodies listed in Table 1. Owing to the latitudinal nonuniformity of rotating convection in spherical shells  (?), the RiNSE formulation cannot be straightforwardly generalized to that setting, but an investigation of the RiNSE with misaligned gravity and rotation axes could provide important insights in this regard. While RiNSE enables simulations at realistic values of ${\rm Ek}$, many challenges remain: the increasing resolution constraint for increasing levels of turbulence cannot be circumvented, and for simplicity our simulations considered stress-free boundaries as well as Prandtl number $\sigma=1$ while ignoring other important ingredients such as magnetic fields. To further test the applicability of our findings to planets, icy moons and stars, future work must also investigate how the regime transitions described in this study are influenced by additional effects, including fluids with Prandtl numbers different from one, compressibility, internal heating  (?), and strong magnetic fields, as well as alternative choices of boundary conditions. The doors to these explorations are now open thanks to the RiNSE formulation. Comparison of the RiNSE numerical results with future laboratory experiments would also be highly desirable.
Materials and Methods
Rescaled Navier-Stokes equations
The convective fluid layer confined by horizontal surfaces separated by a distance $H$ is subjected to rotation around the vertical axis $\boldsymbol{\Omega}=\Omega\mathbf{\widehat{e}}_{z}$ and uniform gravity $-g\mathbf{\widehat{e}}_{z}$. The temperature difference between the bottom and the top plate is $\Delta_{T}>0$. In the Oberbeck-Boussinesq limit, the fluid density varies linearly with temperature $\rho=\rho_{0}(1-\alpha[T-T_{0}])$ near a reference state $(\rho_{0},T_{0})$, with both the thermal expansion coefficient $\alpha$, the fluid kinematic viscosity $\nu$, and the temperature diffusivity $\kappa$ considered constant. The dimensional governing equations for the solenoidal velocity field ($\boldsymbol{\nabla\cdot U}=0$) are:
$\partial_{t}\boldsymbol{U}+\boldsymbol{U\cdot\nabla U}+2\boldsymbol{\Omega}\times\boldsymbol{U}=-\nabla P+\alpha gT\mathbf{\widehat{e}}_{z}+\nu\Delta\boldsymbol{U}$ (3)
$\partial_{t}T+\boldsymbol{U\cdot\nabla}T=\kappa\Delta T.$ (4)
As described in detail in Ref.  (?), the power behind the Rescaled incompressible Navier-Stokes Equations (RiNSE) lies in their use of rotation-influenced characteristic scales for nondimensionalisation. For convenience, we introduce the small parameter $\varepsilon\equiv{\rm Ek}^{1/3}=\left(\nu/2\Omega H^{2}\right)^{1/3}$ and measure lengths in units of the layer depth $H$ in the vertical, but in units of $\ell_{*}=\varepsilon H$ in the horizontal, so that $(X,Y,Z)=(\varepsilon H\widetilde{x},\varepsilon H\widetilde{y},H\widetilde{z})$. In these coordinates, the Laplace operator becomes $\Delta=\left(\varepsilon^{-2}H^{-2}\right)\widetilde{\nabla}_{\varepsilon}^{2}$ where we have introduced the anisotropic diffusion operator: ~$\widetilde{\nabla}_{\varepsilon}^{2}=\partial_{\tilde{x}\tilde{x}}+\partial_{\tilde{y}\tilde{y}}+\varepsilon^{2}\partial_{\tilde{z}\tilde{z}}$. All tildes will be dropped in the following. Time is non-dimensionalized using a diffusive time-scale based on the horizontal scale rather than the layer depth alone: $t_{*}=\ell_{*}^{2}/\nu=\varepsilon^{2}H^{2}/\nu$. Accordingly, the velocity scale is $u_{*}=\ell_{*}/t_{*}=\nu/(\varepsilon H)$, giving the nondimensional velocity $\boldsymbol{u}=\boldsymbol{U}/u_{*}\equiv(u,v,w)$. This isotropic velocity scale, combined with our anisotropic coordinate system, results in horizontally-dominated material derivatives:
$D_{t}^{\perp,\varepsilon}=\partial_{t}+u\partial_{x}+v\partial_{y}+\varepsilon w\partial_{z}=\partial_{t}+\boldsymbol{u_{\perp}\cdot\nabla_{\perp}}+\varepsilon w\partial_{z}.$ (5)
Here the perpendicular symbol $\perp$ denotes the projection upon the horizontal plane: specifically, $\boldsymbol{\nabla}_{\perp}=\left(\partial_{x},\partial_{y},0\right)^{\top}$ and $\boldsymbol{u_{\perp}}=\boldsymbol{u}-w\,\mathbf{\widehat{e}}_{z}$. The incompressibility condition becomes:
$\boldsymbol{\nabla_{\perp}\cdot u_{\perp}}+\varepsilon\partial_{z}w=0.$ (6)
The (dimensional) temperature $T$ is decomposed as the sum of three terms: an unstable linear stratification, a vertical profile $\Theta(z)$, and horizontal fluctuations $\theta(x,y,z)$ with zero horizontal mean:
$T-T_{0}=\Delta_{T}\left[-z+\Theta(z)+\varepsilon\theta(x,y,z)\right].$ (7)
We thus formally embed in our decomposition one hallmark of rapidly rotating convection: horizontal temperature fluctuations are dominated by vertical variations. With these elements, the dimensionless governing equations for the horizontal fluctuations $(\partial_{x},\partial_{y})\neq(0,0)$ of momentum and temperature are found to be:
$D_{t}^{\perp,\varepsilon}\boldsymbol{u}+\varepsilon^{-1}\mathbf{\widehat{e}}_{z}\times\boldsymbol{u}=-\varepsilon^{-1}\nabla p+\frac{\rm\widetilde{Ra}}{\sigma}\theta\,\mathbf{\widehat{e}}_{z}+\widetilde{\nabla}_{\varepsilon}^{2}\boldsymbol{u}$ (8a)
$D_{t}^{\perp,\varepsilon}\theta+w\left(\partial_{z}\Theta-1\right)=\frac{1}{\sigma}\widetilde{\nabla}_{\varepsilon}^{2}\theta,$ (8b)
where $\rm\widetilde{Ra}={\rm Ek}^{4/3}{\rm Ra}$, and the Rayleigh number ${\rm Ra}$ is given by ${\rm Ra}=\alpha gH^{3}\Delta_{T}/(\nu\kappa)$ in terms of the thermal expansion coefficient $\alpha$ of the fluid and gravitational acceleration $g$. By denoting the horizontal average of any quantity using an overline
$\overline{q}^{xy}=\frac{1}{L_{x}L_{y}}\int_{0}^{L_{x}}\mathrm{d}x\int_{0}^{L_{y}}\mathrm{d}y\,q(x,y,z,t)\,,$ (9)
we write separately the equations governing the vertical profiles of horizontal velocity $(\overline{u}^{xy},\overline{v}^{xy})$ and temperature $\Theta$:
$\partial_{t}\overline{u}^{xy}+\varepsilon\partial_{z}\overline{wu}^{xy}-\varepsilon^{-1}\overline{v}^{xy}$ $=\varepsilon^{2}\partial_{zz}\overline{u}^{xy}$ (10a)
$\partial_{t}\overline{v}^{xy}+\varepsilon\partial_{z}\overline{wv}^{xy}+\varepsilon^{-1}\overline{u}^{xy}$ $=\varepsilon^{2}\partial_{zz}\overline{v}^{xy}$ (10b)
$\varepsilon^{-2}\partial_{t}\Theta+\overline{\partial_{z}w\theta}^{xy}$ $=\frac{1}{\sigma}\partial_{zz}\Theta.$ (10c)
Reduced non-hydrostatic quasi-geostrophic equations
In the limit of rapid rotation ${\rm Ek},{\rm Ro}\to 0$, one can derive a reduced set of governing equations for the flow (presented below) in a rigorous and controlled way using asymptotic methods  (?, ?, ?, ?). Here we emphasize, in a quick sketch, how the rescaled equations (8) naturally lead to the reduced equations when the formal limit $\varepsilon\to 0$ is considered. Upon inspection, if the scaling employed for deriving the rescaled equations holds and all dimensionless variables remain $O(1)$, the leading order terms at order $O(\varepsilon^{-1})$ in equation (8a) indicate geostrophic balance:
$\mathbf{\widehat{e}}_{z}\times\boldsymbol{u}=-\nabla p$ (11)
so that the pressure serves as the geostrophic streamfunction:
$\boldsymbol{u}$ =×$=-\boldsymbol{\nabla}\times p\,\mathbf{\widehat{e}}_{z}+w\,\mathbf{\widehat{e}}_{z}+O(\varepsilon)$ (12a)
$=-\partial_{y}p\,\mathbf{\widehat{e}}_{x}+\partial_{x}p\,\mathbf{\widehat{e}}_{y}+w\,\mathbf{\widehat{e}}_{z}+O(\varepsilon).$ (12b)
Substituting expression (12) into the rescaled momentum equation and projecting upon $\mathbf{\widehat{e}}_{z}$, one readily obtains at leading order in $\varepsilon$:
$\partial_{t}w+J[p,w]+\partial_{z}p=\frac{\rm\widetilde{Ra}}{\sigma}\theta+\nabla_{\perp}^{2}w,$ (13a)
[…]
$\partial_{t}\nabla_{\perp}^{2}p+J[p,\nabla_{\perp}^{2}p]-\partial_{z}w=\nabla_{\perp}^{4}p.$ (13b)
Finally, one readily obtains the leading-order contribution to the governing equation for the temperature fluctuation $\theta$ and the temperature mean $\Theta$:
$\partial_{t}\theta+J[p,\theta]+w\left(\partial_{z}\Theta-1\right)$ $=\frac{1}{\sigma}\widetilde{\nabla}_{\varepsilon}^{2}\theta$ (13c)
$\varepsilon^{-2}\partial_{t}\Theta+\overline{\partial_{z}w\theta}^{xy}$ $=\frac{1}{\sigma}\partial_{zz}\Theta.$ (13d)
Equations (13) form the reduced Non-Hydrostatic Quasi-Geostrophic Equations (NHQGE), which were previously derived using a multiple-scale asymptotic expansion  (?).
The Ekman number dependence that remains in the governing equation (13d) for the mean temperature reflects the fact that the timescale $\tau=\varepsilon^{2}t$ associated with the evolution of $\Theta$ is much slower than the convective timescale. Previous studies  (?, ?, ?) have established that, in the limit of horizontally extended domains where a large number of convective cells or plumes contribute to the horizontal spatial averaging, this term can be omitted for the computation of global diagnostics of a statistically stationary state $\partial_{t}\overline{\Theta}\approx 0$. This assumption can be justified based on the following rationale: as the horizontal extent $L_{\perp}=L_{x}=L_{y}$ of the domain increases, so does the number of independent plumes $N_{p}$ that feed back into the temperature profile. The central limit theorem then suggests that, in the ergodic limit (time and ensemble averages behave similarly), the amplitude of temporal fluctuations of the vertical temperature profile $\Theta(z)$ should decrease like $1/\sqrt{N_{p}}\approx 1/L_{\perp}$, justifying the neglect of the term $\partial_{t}\Theta$. In the context of the RiNSE formulation, this assumption has been successfully validated and has proved to shorten the transient that leads to the stationary state  (?).
Numerical simulations
We compute solutions to both the rescaled and the reduced equations using the pseudo-spectral code Coral  (?), a flexible platform for solving partial differential equations in a Cartesian domain. We suppose periodicity along the horizontal directions $(x,y)$ and thus decompose the problem variables using a Fourier basis, whereas a Chebyshev expansion is used in the vertical in order to impose boundary conditions on the bounding surfaces $z=0,1$ using a Galerkin basis recombination. Differential operators are discretized using the quasi-inverse method  (?, ?) to optimize memory usage and accuracy. The resolution was varied from $192$ grid points (corresponding to $128$ alias-free modes) at the lowest values of $\rm\widetilde{Ra}$ in each direction, and up to $768$ grid points ($512$ Chebychev modes) in the vertical direction as well as $588$ grid points ($384$ Fourier modes) in the horizontal directions. The time-marching relies on a family of stable semi-implicit Runge-Kutta schemes  (?), specifically the third-order ARS443 scheme.
For some runs presented here, we have used low amplitude noise as the initial condition, and waited for the convective turbulent flow to settle into a statistically stationary state after a transient period. However, for strongly supercritical flow, this transient may prove much more turbulent than the stationary state itself. Hence, in some cases, we resort to an alternative: instead of low amplitude noise, the initial condition for a given run would be chosen as the solution computed at a neighboring position in parameter space. This procedure facilitates the computation of the transient.
Once the transient has subsided, we compute time averages of different diagnostic quantities reported here. For instance, the barotropic vertical vorticity displayed in Fig. 2 is defined as:
$\overline{\omega_{z}}(x,y)=\int_{0}^{1}\omega_{z}\,\mathrm{d}z,$ (14)
where the vertical vorticity is given by $\omega_{z}=\partial_{x}v-\partial_{y}u$. The Nusselt number is computed in terms of a volume average as ${\rm Nu}(t)-1=\sigma\langle w\theta\rangle_{\rm vol}$, where the volume average is given by
$\langle\cdot\rangle_{\rm vol}=\overline{\overline{(\cdot)}^{xy}}=\frac{1}{L_{x}L_{y}}\int_{0}^{L_{x}}\mathrm{d}x\int_{0}^{L_{y}}\mathrm{d}y\int_{0}^{1}(\cdot)\mathrm{d}z$ (15)
and the small-scale Reynolds number based on the vertical velocity and the horizontal length scale is given, in the nondimensional RiNSE formulation, by ${\rm Ro}=\sqrt{\langle w^{2}\rangle}$, where brackets denote a combined space and time averaging:
$\langle\cdot\rangle=\frac{1}{t_{2}-t_{1}}\int_{t_{1}}^{t_{2}}\mathrm{d}t\frac{1}{L_{x}L_{y}}\int_{0}^{L_{x}}\mathrm{d}x\int_{0}^{L_{y}}\mathrm{d}y\int_{0}^{1}\omega_{z}\,\mathrm{d}z.$ (16)
Acknowledgments
K. J. passed away before this manuscript was finalized. We have attempted to present the results of our collaboration in accordance with his high standards. Any errors or misinterpretations remain our own. This work was supported by the National Science Foundation (Grants DMS-2009319 and DMS-2308338 (KJ), Grants DMS-2009563 and DMS-2308337 (EK)), by the German Research Foundation (DFG Projektnummer: 522026592) and by Agence Nationale de la Recherche (Grant ANR-23-CE30-0016-01). Part of the writing of this manuscript was done while one of the authors (AvK) was a Staff Member at the Woods Hole Oceanographic Institution Geophysical Fluid Dynamics summer program. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California Berkeley (supported by the UC Berkeley Chancellor, Vice Chancellor for Research, and Chief Information Officer). This research also utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University. Data storage for this project was supported by the University of Colorado Boulder PetaLibrary. This project was also granted access to computational resources of TGCC under the allocation A0162A10803 made by GENCI. We thank Daria Holdenried-Chernoff for pointing out useful references on dynamo simulations at low Ekman number.